# Learning to Think: Information-Theoretic Reinforcement Fine-Tuning for LLMs

### v2

@@
Abstract: Large language models (LLMs) excel at complex tasks thanks to advances in their reasoning abilities. However, existing methods overlook the trade-off between reasoning effectiveness and efficiency, often encouraging unnecessarily long reasoning chains and wasting tokens. To address this, we propose Learning to Think (L2T), an information-theoretic reinforcement fine-tuning framework for LLMs to make the models achieve optimal reasoning with fewer tokens. Specifically, L2T treats each query-response interaction as a hierarchical session of multiple episodes and proposes a universal dense process reward, i.e., quantifies the episode-wise information gain in parameters, requiring no extra annotations or task-specific evaluators. We propose a method to quickly estimate this reward based on PAC-Bayes bounds and the Fisher information matrix. Theoretical analyses show that it significantly reduces computational complexity with high estimation accuracy. By immediately rewarding each episode’s contribution and penalizing excessive updates, L2T optimizes the model via reinforcement learning to maximize the use of each episode and achieve effective updates. Empirical results on various reasoning benchmarks and base models demonstrate the advantage of L2T across different tasks, boosting both reasoning effectiveness and efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have progressed from handling basic natural language processing tasks to tackling complex problems, such as writing and maintaining code bases [23, 67, 38, 48], navigating the web and controling devices [66, 54, 17, 4], and acting as personal assistants [21, 30, 22, 7], thanks to the advances in their reasoning abilities. Recent results [24, 19, 56, 43] in LLM reasoning show that scaling test-time compute can substantially improve reasoning capabilities, e.g., [35, 5] demonstrated that generating more tokens during inference yields logarithmic-linear gains. Based on this, a new class of reasoning models [43, 53, 9, 49] has coupled test-time compute scaling with reinforcement learning (RL), achieving state-of-the-art (SOTA) results on various challenging benchmarks [15, 61, 16]. These models employ chain-of-thought (CoT) tokens to guide multi-step reasoning and maintain logical consistency throughout the solution process [50, 52, 58]; by extending and optimizing CoT paths to produce trajectories longer than typical correct solutions, they more thoroughly explore the solution space and thereby boost final answer accuracy [43, 35, 20].
@@
Despite existing methods having demonstrated great performance, they still struggle to balance reasoning effectiveness and efficiency. Specifically, existing approaches typically rely on final outcome rewards for policy optimization, providing no feedback on intermediate reasoning steps. Under such delayed feedback, extending the reasoning chain does not incur any cost, and even a tiny accuracy gain from a large amount of extra reasoning steps is treated as a positive signal [59, 55]. Consequently, the models favor a “one more thought” and continually lengthen their CoTs, resulting in redundant computation and thus reducing reasoning efficiency. Our experiments in Subsection 3.2 further demonstrate this (Figure 1): existing outcome-reward-based RL methods often lead LLMs to consume more than twice the tokens actually needed for the correct answer. Furthermore, by evaluating across different reasoning tasks, we find that this redundancy not only wastes resources but sometimes degrades reasoning effectiveness. For example, on difficult questions (e.g., Tier 4 multi-stage math questions [13]), moderate chain extensions improve coverage of critical steps; whereas on simple tasks (e.g., Tier 1 question “12 + 5”), overly long reasoning chains may reduce overall accuracy. Since real-world tasks vary, no fixed chain length is optimal for all cases. Therefore, designing effective dense process rewards to assess the contribution of each reasoning step is both necessary and valuable. Such rewards help the model to generate tokens that most benefit the answer, ensuring reasoning effectiveness with minimal token budget and efficient learning.
@@
Empirically, across challenging reasoning benchmarks (e.g., AIME, AMC, HumanEval) and base models (e.g., DeepScaleR-1.5B-Preview, DeepSeek-R1-Distill-Qwen-1.5B, DeepSeekR1-Distill-Qwen-7B), L2T consistently achieves comparable performance and stable improvements (Section 5). Compared to standard outcome-reward approaches (e.g., GRPO), it boosts performance by about 3.7% and doubles token efficiency; compared to process-reward baselines (e.g., ReST-MCTS, MRT), it raises accuracy by about 2% and increases efficiency by about 1.3$\times$. In multi-task evaluations, L2T delivers an average accuracy gain of nearly 3% across tasks of varying difficulty and maintains stable improvements under different token budgets. These results demonstrate the advantages of L2T, which effectively balances reasoning effectiveness and efficiency across diverse reasoning scenarios.
@@ 2 Related Work
Complex reasoning has long been recognized as one of the most challenging capabilities for LLMs [23, 67, 19, 56, 43]. To enhance inference, several works [43, 53, 49, 57] have incorporated outcome-reward RL during fine-tuning. RL paradigm mainly extends and optimizes CoT paths based on test-time compute scaling to more thoroughly explore the solution space [35, 5]. However, recent studies [59, 55, 5] show that reasoning length and accuracy are not strictly positively correlated. Excessively long CoTs not only consume undue tokens but also degrade performance (also demonstrated in Subsection 3.2). The wasting token budget reduces the efficiency of unit tokens under a limited budget due to problems such as attention dilution and context truncation, so that the final accuracy may decline [31, 44]. Some methods attempt to reduce reasoning depth via process rewards [63, 36], heuristic scoring [11, 20], or length penalties [57, 1, 33]. These approaches, however, require task-specific evaluators, incurring prohibitive annotation costs with poor cross-task reuse (with more comparison in Appendices C and F). To address these limitations, we propose a universal information-theoretic dense process reward and leverage reinforcement fine-tuning to adaptively recognize reasoning depth across diverse tasks. This design achieves the trade-off between reasoning effectiveness and efficiency without additional annotations or specialized evaluators.
@@
Unlike outcome rewards that assess the final answer, a PRM evaluates the correctness of each intermediate reasoning step [29, 47]. PRMs can be trained via automated supervision without costly human-annotated process labels [34, 39, 60]. Once learned, a PRM can both guide test-time search by allocating additional compute and accelerate exploration in RL using the policy’s own trajectories [43, 35, 60]. Recent works [42, 65, 63, 10, 64] have applied process rewards to fine-tune LLMs with RL, for example, [65] uses relative progress estimation to generate high-quality intermediate supervision labels; [57] derives exploration bonuses from length penalties or an LLM-based judge; and [36] introduces a regret-minimization process reward to optimize test-time compute. However, these methods depend on external annotations or task-specific evaluators, which are expensive to produce and difficult to reuse when task requirements change. In contrast, our approach uses internal model signals, i.e., information gain, as intermediate rewards to optimize the policy without additional interaction data. By constructing a universal dense process reward, our framework applies seamlessly across diverse tasks and promotes efficient, step-wise reasoning.
@@ 3.1 Problem Settings
Our goal is to fine-tune an LLM so that it can answer arbitrary questions $x$ within a fixed token budget $B_{\rm token}$, enforcing efficient reasoning. We treat the LLM as a stochastic policy $\pi_{\theta}(\cdot\mid x)$ parameterized by $\theta$. Each question $x$ is drawn from an underlying distribution $\mathcal{P}_{x}$. Under the token budget $B_{\rm token}$, the model may generate up to $T_{max}$ tokens at test time, and yields an output sequence $z_{0:T}=(z_{0},z_{1},\dots,z_{T})$ for each $x$. The generation process can be formulated as a finite-horizon Markov decision process (MDP): at each time step $t$, the “state” $s_{t}$ consists of the question $x$ together with the partial prefix $(z_{0},\dots,z_{t-1})$, and the “action” $a_{t}$ is the choice of next token $z_{t}$. To guide learning toward correct answers, the MDP is equipped with a reward function $r(s_{t},a_{t})$. It is typically defined as a binary outcome reward [16, 20] to determine whether the generated answer is correct.
According to existing RL-based fine-tuning paradigm [41, 43, 16], we are given a training dataset $\mathcal{D}_{\rm train}=\{(x_{i},y^{*}_{i})\}_{i=1}^{N}$, where each $y^{*}_{i}$ is an oracle reasoning trace that leads to the correct answer. During fine-tuning, we use these traces both to calculate the reward and to guide learning. For each question $x_{i}$, we first generate candidate token sequences $z_{0:T}\sim\pi_{\theta}(\cdot\mid x_{i})$. Then, we compute the reward $r(x_{i},z_{0:T})$, which equals 1 if $z_{0:T}$ matches the oracle trace $y^{*}_{i}$. By maximizing the empirical sum of these rewards with the constraint of test-time budget $B_{\rm token}$, we update $\pi_{\theta}$. The objective is:
@@
Figure 1: Results of DeepScaleR-1.5B-Preview across different tasks on Omni-MATH. We partition the generated reasoning chain into episodes, measuring accuracy $\mathrm{Acc}(k)$ and average token consumption $\overline{T}(k)$ at different episode depths. More details and results are shown in Appendix F.1.
@@ 3.2 Empirical Evidence
To ensure reasoning effectiveness, existing methods [43, 9, 53, 15, 49] mainly leverage test-time compute scaling to lengthen reasoning chains beyond what is minimally required for a correct answer, thereby more thoroughly exploring the solution space and improving final answer accuracy. These approaches typically use a sparse outcome reward for policy optimization, without feedback on intermediate reasoning steps [59, 55]. Using this framework, extending the chain carries no penalty, and even minimal accuracy gains from extra steps yield a positive signal. Thus, models resort to consuming additional tokens to secure the correctness of the final answer [31, 35]. However, under the fixed token budget $B_{token}$ considered in this paper, this extension may deplete the budget prematurely, undermining efficient reasoning in subsequent steps. To validate this argument, in this subsection, we conduct a series of experiments to analyze the token-utilization abilities of existing methods.
@@
Specifically, in this experiment, we use the Omni-MATH benchmark [13] for evaluation, which comprises 4,428 questions across more than 33 domains, including algebra, discrete mathematics, geometry, number theory, etc. The questions are organized by human experts into four difficulty tiers (Tier 1-4) to assess model performance on tasks of varying complexity. We evaluate two base models [16], i.e., DeepScaleR-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B, running on the A100 GPU clusters with greedy decoding (temperature = 0) and a maximum generation length of 16,384 tokens. To probe different reasoning depths, we wrap each logical step in the prompt with ‘<think>’…‘</think>’ tags, segmenting the generated chain into up to $K=30$ episodes. For each test question and each truncation point $k=1,\dots,30$, we force-stop generation at the $k$-th ‘</think>’ tag, append the prompt “Please output the final answer directly based on the above steps”, and then use Omni-Judge to determine correctness. This yields the accuracy $\mathrm{Acc}(k)$ at episode depth $k$ and the average token consumption $\overline{T}(k)$ from the initial prompt to the truncation. To mitigate the impact of randomness, we introduce a $\mathrm{maj@4}(k)$ baseline: for each truncated context we sample four continuations, take the majority-vote result as the final answer, and record its accuracy $\mathrm{maj@4}(k)$ and token cost. By plotting “accuracy vs. episode number” and “token consumption vs. episode number”, we can visualize how performance evolves with reasoning depth across tasks of differing difficulty and compare sequential generation against majority-vote under the same compute budget.
@@
From Figure 1, we can observe that: (i) Existing methods may fail to use test-time compute budgets efficiently, leading to wasted resources: both models have on average used more than twice the minimum tokens required. For example, $k=16$ achieves accuracy comparable to or exceeding sequential generation at $k=24$ with fewer tokens. (ii) The additional episodes add no new information and instead degrade performance due to context redundancy: for both models, $\mathrm{Acc}(k)$ peaks around $k\approx 16-20$ and then declines as $k$ increases. (iii) The questions of different difficulty tiers prefer different chain lengths: Tier 4 questions tend to benefit from longer chains, whereas Tier 1 questions can achieve correct results with short chains, where excessive reasoning depth may causes a marked accuracy drop (e.g., falls by over 5% at $k=20$). These findings underscore the limitations of existing methods, which ignore the balance between reasoning effectiveness and efficiency.

### v3

@@
Abstract
Large language models (LLMs) excel at complex tasks thanks to advances in their reasoning abilities. However, existing methods overlook the trade-off between reasoning effectiveness and efficiency, often encouraging unnecessarily long reasoning chains and wasting tokens. To address this, we propose Learning to Think (L2T)  [^1] Project page: https://wangjingyao07.github.io/L2T.github.io/, an information-theoretic reinforcement fine-tuning framework for LLMs to make the models achieve optimal reasoning with fewer tokens. Specifically, L2T treats each query-response interaction as a hierarchical session of multiple episodes and proposes a universal dense process reward, i.e., quantifies the episode-wise information gain in parameters, requiring no extra annotations or task-specific evaluators. We propose a method to quickly estimate this reward based on PAC-Bayes bounds and the Fisher information matrix. Theoretical analyses show that it significantly reduces computational complexity with high estimation accuracy. By immediately rewarding each episode’s contribution and penalizing excessive updates, L2T optimizes the model via reinforcement learning to maximize the use of each episode and achieve effective updates. Empirical results on various reasoning benchmarks and base models demonstrate the advantage of L2T across different tasks, boosting both reasoning effectiveness and efficiency.
@@ 1 Introduction
Large Language Models (LLMs) have progressed from handling basic natural language processing tasks to tackling complex problems, such as writing and maintaining code bases jimenez2024swebenchlanguagemodelsresolve ; zhou2025sweetrltrainingmultiturnllm ; sadik2025benchmarkingllmcodesmells ; wang2025insightsverificationtrainingverilog , navigating the web and controling devices zhou2024webarenarealisticwebenvironment ; xu2024aguvisunifiedpurevision ; gur2024realworldwebagentplanninglong ; bai2024digirltraininginthewilddevicecontrol , and acting as personal assistants jiang2025knowmerespondme ; liu2025surveypersonalizedlargelanguage ; jiang2024delegationdesigningidealagentic ; chaudhuri2025tripcraftbenchmarkspatiotemporallyfine , thanks to the advances in their reasoning abilities. Recent results kaplan2020scalinglawsneurallanguage ; hoffmann2022trainingcomputeoptimallargelanguage ; yao2023treethoughtsdeliberateproblem ; snell2024scalingllmtesttimecompute in LLM reasoning show that scaling test‑time compute can substantially improve reasoning capabilities, e.g., muennighoff2025s1 ; ballon2025relationship demonstrated that generating more tokens during inference yields logarithmic‑linear gains. Based on this, a new class of reasoning models snell2024scalingllmtesttimecompute ; wu2025phdknowledgerequiredreasoning ; chen2025think23overthinkingo1like ; wang2025thoughtsplaceunderthinkingo1like has coupled test‑time compute scaling with reinforcement learning (RL), achieving state‑of‑the‑art (SOTA) results on various challenging benchmarks guan2025rstarmathsmallllmsmaster ; yuenyong2025openthaigpt16r1thaicentric ; deepseekai2025deepseekr1incentivizingreasoningcapability . These models employ chain‑of‑thought (CoT) tokens to guide multi‑step reasoning and maintain logical consistency throughout the solution process wei2023chainofthoughtpromptingelicitsreasoning ; wu2025inferencescalinglawsempirical ; ye2025limoreasoning ; by extending and optimizing CoT paths to produce trajectories longer than typical correct solutions, they more thoroughly explore the solution space and thereby boost final answer accuracy snell2024scalingllmtesttimecompute ; muennighoff2025s1 ; ji2025testtimecomputesystem1thinking .
@@
Despite existing methods having demonstrated great performance, they still struggle to balance reasoning effectiveness and efficiency. Specifically, existing approaches typically rely on final outcome rewards for policy optimization, providing no feedback on intermediate reasoning steps. Under such delayed feedback, extending the reasoning chain does not incur any cost, and even a tiny accuracy gain from a large amount of extra reasoning steps is treated as a positive signal yeo2025demystifyinglongchainofthoughtreasoning ; yang2025thinkneedselfadaptivechainofthought . Consequently, the models favor a “one more thought” and continually lengthen their CoTs, resulting in redundant computation and thus reducing reasoning efficiency. Our experiments in Subsection 3.2 further demonstrate this (Figure 1): existing outcome-reward-based RL methods often lead LLMs to consume more than twice the tokens actually needed for the correct answer. Furthermore, by evaluating across different reasoning tasks, we find that this redundancy not only wastes resources but sometimes degrades reasoning effectiveness. For example, on difficult questions (e.g., Tier 4 multi-stage math questions gao2024omnimathuniversalolympiadlevel ), moderate chain extensions improve coverage of critical steps; whereas on simple tasks (e.g., Tier 1 question “12 + 5”), overly long reasoning chains may reduce overall accuracy. Since real-world tasks vary, no fixed chain length is optimal for all cases. Therefore, designing effective dense process rewards to assess the contribution of each reasoning step is both necessary and valuable. Such rewards help the model to generate tokens that most benefit the answer, ensuring reasoning effectiveness with minimal token budget and efficient learning.
@@
[…]
@@ 2 Related Work
Complex reasoning has long been recognized as one of the most challenging capabilities for LLMs jimenez2024swebenchlanguagemodelsresolve ; zhou2025sweetrltrainingmultiturnllm ; hoffmann2022trainingcomputeoptimallargelanguage ; yao2023treethoughtsdeliberateproblem ; snell2024scalingllmtesttimecompute . To enhance inference, several works snell2024scalingllmtesttimecompute ; wu2025phdknowledgerequiredreasoning ; wang2025thoughtsplaceunderthinkingo1like ; ye2025emergence have incorporated outcome-reward RL during fine-tuning. RL paradigm mainly extends and optimizes CoT paths based on test-time compute scaling to more thoroughly explore the solution space muennighoff2025s1 ; ballon2025relationship . However, recent studies yeo2025demystifyinglongchainofthoughtreasoning ; chen2025think23overthinkingo1like ; yang2025thinkneedselfadaptivechainofthought ; ballon2025relationship show that reasoning length and accuracy are not strictly positively correlated. Excessively long CoTs not only consume undue tokens but also degrade performance (also demonstrated in Subsection 3.2). The wasting token budget reduces the efficiency of unit tokens under a limited budget due to problems such as attention dilution and context truncation, so that the final accuracy may decline liu2024mindstepbystep ; stechly2025chainthoughtlessnessanalysiscot . Some methods attempt to reduce reasoning depth via process rewards zhang2024entropyregularizedprocessrewardmodel ; qu2025optimizing , heuristic scoring feng2024alphazeroliketreesearchguidelarge ; ji2025testtimecomputesystem1thinking , or length penalties ye2025emergence ; aggarwal2025l1 ; luo2025o1 . These approaches, however, require task-specific evaluators, incurring prohibitive annotation costs with poor cross-task reuse (with more comparison in Appendices D and G). To address these limitations, we propose a universal information-theoretic dense process reward and leverage reinforcement fine-tuning to adaptively recognize reasoning depth across diverse tasks. This design achieves the trade-off between reasoning effectiveness and efficiency without additional annotations or specialized evaluators.
@@
Unlike outcome rewards that assess the final answer, a PRM evaluates the correctness of each intermediate reasoning step lightman2023let ; uesato2022solving . PRMs can be trained via automated supervision without costly human-annotated process labels luo2024improve ; setlur2024rewarding ; yuan2024free . Once learned, a PRM can both guide test-time search by allocating additional compute and accelerate exploration in RL using the policy’s own trajectories snell2024scalingllmtesttimecompute ; muennighoff2025s1 ; yuan2024free . Recent works she2025rprmreasoningdrivenprocessreward ; zhao2025genprmscalingtesttimecompute ; zhang2024entropyregularizedprocessrewardmodel ; cheng2025stopsummationminformcredit ; zhang2025lessonsdevelopingprocessreward have applied process rewards to fine-tune LLMs with RL, for example, zhao2025genprmscalingtesttimecompute uses relative progress estimation to generate high-quality intermediate supervision labels; ye2025emergence derives exploration bonuses from length penalties or an LLM-based judge; and qu2025optimizing introduces a regret-minimization process reward to optimize test-time compute. However, these methods depend on external annotations or task-specific evaluators, which are expensive to produce and difficult to reuse when task requirements change. In contrast, our approach uses internal model signals, i.e., information gain, as intermediate rewards to optimize the policy without additional interaction data. By constructing a universal dense process reward, our framework applies seamlessly across diverse tasks and promotes efficient, step-wise reasoning.
@@ 3.1 Problem Settings
Our goal is to fine-tune an LLM so that it can answer arbitrary questions $x$ within a fixed token budget $B_{\rm token}$, enforcing efficient reasoning. We treat the LLM as a stochastic policy $\pi_{\theta}(\cdot\mid x)$ parameterized by $\theta$. Each question $x$ is drawn from an underlying distribution $\mathcal{P}_{x}$. Under the token budget $B_{\rm token}$, the model may generate up to $T_{max}$ tokens at test time, and yields an output sequence $z_{0:T}=(z_{0},z_{1},\dots,z_{T})$ for each $x$. The generation process can be formulated as a finite-horizon Markov decision process (MDP): at each time step $t$, the “state” $s_{t}$ consists of the question $x$ together with the partial prefix $(z_{0},\dots,z_{t-1})$, and the “action” $a_{t}$ is the choice of next token $z_{t}$. To guide learning toward correct answers, the MDP is equipped with a reward function $r(s_{t},a_{t})$. It is typically defined as a binary outcome reward deepseekai2025deepseekr1incentivizingreasoningcapability ; ji2025testtimecomputesystem1thinking to determine whether the generated answer is correct.
According to existing RL-based fine-tuning paradigm shao2024deepseekmathpushinglimitsmathematical ; snell2024scalingllmtesttimecompute ; deepseekai2025deepseekr1incentivizingreasoningcapability , we are given a training dataset $\mathcal{D}_{\rm train}=\{(x_{i},y^{*}_{i})\}_{i=1}^{N}$, where each $y^{*}_{i}$ is an oracle reasoning trace that leads to the correct answer. During fine-tuning, we use these traces both to calculate the reward and to guide learning. For each question $x_{i}$, we first generate candidate token sequences $z_{0:T}\sim\pi_{\theta}(\cdot\mid x_{i})$. Then, we compute the reward $r(x_{i},z_{0:T})$, which equals 1 if $z_{0:T}$ matches the oracle trace $y^{*}_{i}$. By maximizing the empirical sum of these rewards with the constraint of test-time budget $B_{\rm token}$, we update $\pi_{\theta}$. The objective is:
@@
Figure 1: Results of DeepScaleR-1.5B-Preview across different tasks on Omni-MATH. We partition the generated reasoning chain into episodes, measuring accuracy $\mathrm{Acc}(k)$ and average token consumption $\overline{T}(k)$ at different episode depths. More details and results are shown in Appendix G.1.
@@ 3.2 Empirical Evidence
To ensure reasoning effectiveness, existing methods snell2024scalingllmtesttimecompute ; chen2025think23overthinkingo1like ; wu2025phdknowledgerequiredreasoning ; guan2025rstarmathsmallllmsmaster ; wang2025thoughtsplaceunderthinkingo1like mainly leverage test-time compute scaling to lengthen reasoning chains beyond what is minimally required for a correct answer, thereby more thoroughly exploring the solution space and improving final answer accuracy. These approaches typically use a sparse outcome reward for policy optimization, without feedback on intermediate reasoning steps yeo2025demystifyinglongchainofthoughtreasoning ; yang2025thinkneedselfadaptivechainofthought . Using this framework, extending the chain carries no penalty, and even minimal accuracy gains from extra steps yield a positive signal. Thus, models resort to consuming additional tokens to secure the correctness of the final answer liu2024mindstepbystep ; muennighoff2025s1 . However, under the fixed token budget $B_{token}$ considered in this paper, this extension may deplete the budget prematurely, undermining efficient reasoning in subsequent steps. To validate this argument, in this subsection, we conduct a series of experiments to analyze the token-utilization abilities of existing methods.
@@
Specifically, in this experiment, we use the Omni-MATH benchmark gao2024omnimathuniversalolympiadlevel for evaluation, which comprises 4,428 questions across more than 33 domains, including algebra, discrete mathematics, geometry, number theory, etc. The questions are organized by human experts into four difficulty tiers (Tier 1-4) to assess model performance on tasks of varying complexity. We evaluate two base models deepseekai2025deepseekr1incentivizingreasoningcapability , i.e., DeepScaleR-1.5B-Preview and DeepSeek-R1-Distill-Qwen-1.5B, running on the A100 GPU clusters with greedy decoding (temperature = 0) and a maximum generation length of 16,384 tokens. To probe different reasoning depths, we configure the prompt to wrap each logical step with ‘<think>’…‘</think>’ tags, segmenting the generated chain into up to $K=30$ episodes. For each test question and each truncation point $k=1,\dots,30$, we force-stop generation at the $k$-th ‘</think>’ tag, append the prompt “Please output the final answer directly based on the above steps”, and then use Omni-Judge to determine correctness. This yields the accuracy $\mathrm{Acc}(k)$ at episode depth $k$ and the average token consumption $\overline{T}(k)$ from the initial prompt to the truncation. Considering the impact of randomness, we introduce a $\mathrm{maj@4}(k)$ baseline: for each truncated context, we sample four continuations, take the majority-vote result as the final answer, and record its accuracy $\mathrm{maj@4}(k)$ and token cost. By plotting “accuracy vs. episode number” and “token consumption vs. episode number”, we can visualize how performance evolves with reasoning depth across tasks of differing difficulty and compare sequential generation against majority-vote under the same compute budget.
@@
From Figures 1 and 6, we can observe that: (i) Existing methods may fail to use test-time compute budgets efficiently, leading to wasted resources: both models have on average used more than twice the minimum tokens required. For example, $k=16$ achieves accuracy comparable to or exceeding sequential generation at $k=24$ with fewer tokens. (ii) The additional episodes add no new information and instead degrade performance due to context redundancy: for both models, $\mathrm{Acc}(k)$ peaks around $k\approx 16-20$ and then declines as $k$ increases. (iii) The questions of different difficulty tiers prefer different chain lengths: Tier 4 questions tend to benefit from longer chains, whereas Tier 1 questions can achieve correct results with short chains, where excessive reasoning depth may causes a marked accuracy drop (e.g., falls by over 5% at $k=20$). These findings underscore the limitations of existing methods, which ignore the balance between reasoning effectiveness and efficiency.